Due to a kink in the bellow on the upstream beam line I decided to replace the beam line and restart the pump down.
The bellows is now straight and the vacuum is on the turbos and coming down.
The alignment went well and we are within ½ mm (thanks alignment group) of where we started accept for Z.
Due to the sequence of the first assembly the shrinkage of the transition O rings was not taken into account so this time we are 1mm farther downstream in Z because we did this survey under vacuum. The data will be made available as soon as it is processed.